transform the quadratic function y= -2x^2-12x-10 into vertex form. find the vertex, x and y intercepts.
Answer by ewatrrr(24785)

y= -2x^2-12x-10 P(0, -10) y-intercept
y = -2(x+3)^2 + 18 - 10
y = -2(x+3)^2 + 8
V(-3,8)
......
0 = -2(x+3)^2 + 8
4 =(x+3)^2
± 2 = x+3
-3 ± 2 = x
P(-5,0) and P(-1,0) x-intercepts